OpenStudy (anonymous):

use rational exponents to write ^2sqrt5 x ^3sqrt8 as a single radical expression

OpenStudy (anonymous):

Do you mean : $$ \large \huge\sqrt[2]{5} \times \sqrt[3]{8} = ?$$

OpenStudy (anonymous):

If yes then the answer is $$2 \sqrt{5}$$

OpenStudy (anonymous):

NOTE : $$ \sqrt[2]{5} = \sqrt{5}$$